### v1

@@ 3.2.1 Implementation Details
The generalizable INR for OCT interpolation consists of a six-layer MLP using complex Garbor wavelet activation functions (saragadam2023wire) plus two separate linear layers as decoder heads for reconstruction output and segmentation output . All hidden layers have 128 nodes, and the reconstruction head outputs scalar grayscale values, while the segmentation head generates output values for one-hot encoded class labels, where is the number of classes considered in the respective task. That is, for the segmentation of retinal layers, twelve labels corresponding to the different tissue layers – including the choroid – are predicted, and for the segmentation of pathological images, three labels corresponding to the background, retina, and pathologies are predicted. The reconstruction output layer uses sigmoid activation, whereas the segmentation head uses softmax activation. All hidden layers use residual connections, and to enhance focus on input coordinates as well as the additional imaging modality, we pass these inputs repeatedly to the hidden layers. The length of the latent vectors is set to 128 empirically. Network training is performed for 1500 epochs, where one epoch corresponds to a run over all B-scans of the training images. We use an exponential learning rate decay of 0.99 and an initial learning rate of 10-4. Early stopping is enabled after 100 epochs using a patience of 20 epochs and a minimum loss improvement of 10-5. The generalizable INR was trained for up to 3 days on a single NVIDIA GeForce GTX 1080 Ti.
@@ 3.3 Registration and Atlas Generation
The learnable atlas is a key component to enable resolution independence. It builds upon the idea proposed by grossbrohmer2024sina of representing the atlas using an MLP with adjustable parameters . As such, the atlas is trained using warped input coordinates to predict the most similar subject representation. In this work, we extend this approach by incorporating a combined representation of intensity values and segmentation labels, similar to our previously proposed interpolation method. Thus, we introduce an INR , which takes warped 3D coordinates to reconstruct the intensity value of the OCT image and the class label of the retinal layers. Since we aim to construct a single atlas volume that represents the whole dataset, this INR is (different from our previous approach) non-generalizable. This continuous representation of the template enables the intensity and segmentation values to be sampled instantly, without the need for explicit interpolation in the registration process.
The displacement model is a generalizable INR, denoted as with trainable parameters . In contrast to the proposed method for interpolation, the displacement model uses a hypernetwork to modulate its interim outputs, similar to dannecker2024cina and dupont2022data. To do so, we use the hypernetwork (depicted in purple in Fig. 3) – consisting of one linear layer for each layer of the INR (depicted in turquoise in Fig. 3) – to predict individual scaling and shifting vectors based on the latent prior input . Let be the hidden size of the INR. The individual scaling vector and shifting vector are the result of the linear mapping with weight matrix , and bias . By modulating with scaling vector and shifting vector , the output of a linear layer with input , weight matrix , bias and activation function becomes .
The resulting latent code-driven architecture is then used to represent a continuous displacement field for each subject as already introduced for classic INRs by wolterink2022implicit. Thus, this model gets the original 3D input coordinates and a latent prior . For each coordinate , we thereby obtain a subject-specific displacement
@@
In contrast to the architecture of dannecker2024cina, we separate the parts and of the latent prior that predict and and use two normal distributions and to initialize them. With and the resulting generalizable INR corresponds to a classic INR. This initialization therefore reduces the disturbance from a non-trained latent prior right after initialization.
Training and inference for atlas generation work analogously to those for interpolation. Assembled to one framework, a subset of coordinates of an OCT volume is used to compute the reconstruction and alignment error as
@@ 3.3.1 Implementation Details
The atlas representation is a four-layer MLP with two additional separate layers for reconstruction and segmentation . Each hidden layer has 256 nodes and uses a sinusoidal activation function as proposed by sitzmann2020implicit. The prediction heads are based on sigmoid activation for the image intensity values and Softmax activation to predict the class values as one-hot encoding. As an extension to the interpolation task, the class labels were redefined to ten retinal layers by merging very thin layers and discarding the choroid label. While grossbrohmer2024sina suggest to use a warm-start procedure by increasing the learning rate, we propose to pre-train the atlas INR for 500 epochs to represent the voxel-wise median of the dataset, resulting in a less blurred initialization. The generalizable INR used to model the deformation consists of four layers with 128 nodes and uses sine activations throughout, including a single sine activation in the final layer to predict the three-dimensional displacement vector. The hypernetwork’s layers are three separate linear layers with 256 nodes each. The model weights are initialized by sampling from a uniform distribution for small and smooth initial deformations. The vector length of 128 for the latent space is chosen empirically. All components – the deformation model, the atlas model, and the latent priors – are trained simultaneously for 2,000 iterations using the Adam optimizer with learning rates of 10-5, 10-6, and 10-5, respectively. The optimizer is combined with an exponential learning rate scheduler with a multiplication factor of 0.999. The batch size is set to one subject, with 916,608 coordinates each. This number corresponds to an equidistant downsampling to a sub-volume containing 2316264 voxels. This volume is randomly selected from all shifted variations of downsampled sub-volumes. Training was conducted on two NVIDIA RTX A5000 and completed in approximately 8 hours. During inference, one latent prior is trained for each subject for 100 epochs with a learning rate of 510-4 and the same exponential learning rate scheduler as for training.
@@ 4.1 Anisotropic Interpolation
The interpolation capabilities of the proposed INR are first assessed using the healthy dataset. To do so, the network is trained with 80 OCT volumes and SLO images of 40 volunteers using only 16 B-scans per volume as input to the INR, resulting in even higher inter-slice distances than for typical clinical images. The input B-scans are sampled equidistantly from the entire depth of the image volumes, enabling evaluation on intermediate B-scans. For comparison, we use simple linear interpolation, registration-based interpolation (ehrhardt2007structure) and non-generalizable INRs adapted to single OCT images. To segment intermediate B-scans for the interpolation methods, we proceed as follows: For the linear interpolation, we interpolate the label images linearly and round the resulting continuous values to integers. For the registration-based method, we use the deformation fields resulting from the B-scan interpolation and apply them to the label images using nearest neighbor interpolation. The non-generalizable INRs (SingleINRs) are implemented with a slightly changed architecture, since we observed improved performance for the modified architecture. The SingleINR, thus, consists of three linear layers of size 512 with sinusoidal activation functions (sitzmann2020implicit). Furthermore, to get the SingleINR to interpolate between the given B-scans instead of overfitting to the given positions, we intentionally use smaller B-scan distances than the real ones so that the input coordinates lie approximately on an isotropic grid.
@@
Since the proposed generalizable INR (GenINR) is the only one of the considered methods that can be used to interpolate and segment unseen test cases, we additionally evaluate the GenINR for the 20 unseen cases from the healthy dataset. For these cases, new latent priors are learned, while the INR weights remain frozen. We repeat this experiment for the pathological data, using 120 OCT and FAF images from 14 patients for training and 26 images from five patients for testing. Here, we use every second B-scan for training, i.e., 13 B-scans per volume, and evaluate performance for the remaining twelve B-scans. Results for the healthy dataset can be found in Tab. 1 and in Tab. 2 for the CSCR dataset. For both datasets, several image similarity metrics, including mean absolute error (MAE), peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM), and learned perceptual image patch similarity (LPIPS), are reported to assess the image reconstruction and interpolation capability of the considered methods. For a consistent evaluation of both proposed approaches, we used the redefined class labels in which the choroid is discarded and very thin layers are merged. To evaluate the segmentation performance, the Dice score averaged over the 9 retinal layers, as well as the average symmetric surface distance (ASSD) and Hausdorff distance (HD) of the layer borders, are reported for the healthy dataset. For the CSCR dataset, we evaluate the segmentation performance with Dice scores of the retina () and of the three disease-related fluid accumulations combined into one binary pathology label ().
@@
The quantitative results for the healthy dataset show that our proposed GenINR achieves the best results for three out of four image similarity metrics. Only for the LPIPS metric, the GenINRs underperform, while the simple interpolation methods (linear and registration-based) achieve the highest values. Simple interpolation hardly changes the contrast and noise level, whereas the GenINR interpolation results look comparably smooth, missing the noisy and high-frequency components of the images (cf. Fig. 4), which might explain the LPIPS results. Regarding image segmentation, the proposed method performs better than the instance-based methods, indicating a better interpolation of the retinal shape achieved by generalization using population-based training. Especially in the fovea region, where large shape differences can be observed for neighboring B-scans, the GenINR is the only method that can meaningfully represent the shape, as shown in the examples on the left side in Fig. 4. The SingleINR, in turn, shows a strong overfitting on the slices used for training, hindering it from interpolating the retinal shape correctly. This overfitting also prevents the SingleINR from benefiting from the additional SLO information; in fact, the additional information, which is different at the interpolated positions than at those used for training, causes SingleINR to fail. The GenINR in turn is enabled to reconstruct blood vessels as shown in the examples on the right in Fig. 4. The localization information provided by the en-face SLO enables the correct positioning of blood vessels and the reconstruction of blood vessel shadows in the outer retina. Reconstruction in the fovea region, where fewer blood vessels are located, appears to be more challenging, as no reconstructions of vessels or vessel shadows are notable. Interpolation artifacts from neighboring B-scans in the form of misplaced shadows, as observed in competing methods, do not occur for the GenINR – all observed vessel shadows are correctly placed. Furthermore, the GenINR enables the interpolation and segmentation of unseen test cases, which is not possible with the instance-specific methods. Although a decline in performance for the test cases compared to the training images can be observed, the GenINR still manages to reconstruct the OCT volumes reliably and to represent the retinal layer boundaries with an average ASSD of .
@@
Similar results can be observed for the CSCR data. Again, the GenINR demonstrates better performance than the baseline methods both in terms of image reconstruction and segmentation, except for the LPIPS metric. As for the healthy data, the simple linear interpolation delivers the best results for this metric. As observed before, the retinal shape is best aligned with the generalizable INR (average Dice score of 98.98 % for the retina) and also the pathologies are better interpolated. For the test data, all metrics considered show similar performance to the training data, except for the Dice score of pathologies. Without FAF integration, an average Dice score of 16.84 % is achieved for the interpolated slices of the test images and improved slightly to 17.59 % with the additional FAF modality. Although there is no known clear relationship between FAF and OCT appearance, the INR seems to benefit from the additional modality, indicating a direct correlation between FAF and OCT. The reduced segmentation performance for the test cases indicates that the highly variable appearance of retinal pathologies in CSCR cannot yet be completely covered by the individualization through the latent codes. Some exemplary interpolation results for the CSCR data are shown in Fig. 5. Here, it can be seen that the retinal layers are reconstructed well for B-scans without severe pathological changes. However, no reconstruction of blood vessels can be observed, which we attribute to the distorting effect of pathologies on training. Also, in pathological areas the retinal layers appear more blurred compared to the regions unaffected by fluids. Still, the shape of both the retina and the pathologies is accurately reproduced by the GenINR.
@@
Across all experiments, the runtime for each subject was below 1 second for linear interpolation, around 19 seconds for registration-based interpolation, approximately 3 minutes for the GenINR (inference adaptation to a new subject), and about 5 minutes for the SingleINR.
@@ 4.2 Atlas Registration
This section assesses the usage of generalizable INRs for image registration tasks – evaluating how effectively they capture structural consistency within a population. Registration performance and atlas generation capabilities are evaluated using the healthy dataset and compared with two different approaches. First, we compared our GenINR framework with the learnable implicit neural atlas approach of grossbrohmer2024sina (SINA), which uses B-spline control points for registration. For the purpose of supervised segmentation, SINA is extended by a second INR that acts as a label atlas, which introduces weak supervision into the atlas generation process. In addition, we modified our proposed method with an non-generalizable INR (SingleINR). Since SingleINRs are trained on individual instances, joint optimization of atlas and deformation presents new challenges. Thus, we used the computed voxel-wise median as a fixed atlas that is registered using the inter-subject approach of wolterink2022implicit. The SingleINR consists of a SIREN with four hidden linear layers of size 128 sitzmann2020implicit. As a reference, we include the initial alignment to the atlas (GenINR), and the alignment after affine registration (GenINR).
@@
Intensity similarity is evaluated using the SSIM, while anatomical alignment is evaluated using the Dice and the ASSD. Deformation’s integrity and plausibility are evaluated using the percentage of negative Jacobian determinants () and the size of the displacements, as measured by the L1 norm (). All similarity metrics are evaluated on single B-scans and averaged over each volume. The methods are trained using the healthy dataset with 64 equidistantly sampled B-scans. The experiments are performed using five-fold cross-validation with 80 OCT and segmentation volumes for training and 20 volumes for inference. We conducted four experiments to evaluate the performance of our proposed registration framework. We 1) evaluate the atlas generation capabilities, 2) investigate the registration performance, 3) show that the framework can be used independently of the inference resolution, and 4) perform an ablation study by replacing the implicit atlas representation by an explicit discrete parameter representation.
@@
Our proposed registration framework creates a well-aligned mean representation as indicated by GenINR in Tab. 3. It achieves scores of , , and for ASSD, Dice, and SSIM, respectively. These findings are supported by the qualitative results shown in Fig. 7. We consider an atlas to be good when it is sharp – especially on the layer edges, artifact-free, and anatomically plausible. Our framework shows qualitative sharper and more precise layer separation as well as hardly any artifacts compared to the median representation or SINA. Moreover, SINA exhibits artifacts at the center of the fovea and, in general, a blurrier representation. While the median atlas already shows minimal artifacts and reasonable layer separation, our method achieves higher sharpness and better separation in the upper retinal layers. Although both generated atlases show a consistent alignment between class labels and intensity representations, our method yields more anatomically plausible results, particularly in the innermost retinal layers, with a continuous representation of the RNFL across the retina.
@@
Focusing on the results of the atlas-to-subject registration experiment in Tab. 3, our approach achieved the best results by about , , and for ASSD, Dice, and SSIM, respectively. Significant improvements are shown in Fig. 6 for GenINR over the initial alignment, affine alignment, and the two approaches that were compared: SINA and SingleINR. Meanwhile, the time taken for inference execution is reduced by around four minutes per subject compared to SINA and around one minute per subject compared to the SingleINR – resulting in faster processing at around per subject. However, SINA presents a significant deterioration in the segmentation metrics compared to the base alignment (GenINR), indicating an insufficient number of control points of the interpolated displacement field. SINA and GenINR produce smooth and fold-free deformation without the need for additional regularization. SingleINR exhibits a slight percentage of folding. Fig 8 presents the anatomical alignment results using the separated ASSD metric for all layers. Due to pre-processing, the bottom layers (RPE, OPR & SVS, and IS/OS) are already well aligned initially and challenge the methods to generate a proper transition from small to larger displacements. Despite this, our framework achieved a uniform performance distributed across all layers, whereas the SingleINR struggled with the upper middle layers GCL, IPL, and INL.
@@
We conducted an ablation study to illustrate the performance gains of the implicit atlas representation. To do so, we replaced the INR with two parameter images for intensities and class labels, which are then transformed using the deformation field and linear interpolation to be compared with the target volume. This explicit atlas representation scales with the size of the three dimensions , which limited the experiment to a resolution of 16 B-scans. The results are shown in Fig. 4 and support the advantage of implicit atlas representation by even improving alignment performance across all segmentation-based metrics. It should be noted that the training time refers to the latent prior training during inference.

### v3

@@ 3.2.1 Implementation Details
The generalizable INR for OCT interpolation consists of a six-layer MLP using complex Garbor wavelet activation functions (Saragadam et al., 2023) plus two separate linear layers as decoder heads for reconstruction output and segmentation output . All hidden layers have 128 nodes, and the reconstruction head outputs scalar grayscale values, while the segmentation head generates output values for one-hot encoded class labels, where is the number of classes considered in the respective task. That is, for the segmentation of retinal layers, twelve labels corresponding to the different tissue layers – including the choroid – are predicted, and for the segmentation of pathological images, three labels corresponding to the background, retina, and pathologies are predicted. The reconstruction output layer uses sigmoid activation, whereas the segmentation head uses softmax activation. All hidden layers use residual connections, and to enhance focus on input coordinates as well as the additional imaging modality, we pass these inputs repeatedly to the hidden layers. The length of the latent vectors is set to 128 empirically. Network training is performed for 1,500 epochs, where one epoch corresponds to a run over all B-scans of the training images. We use an exponential learning rate decay of 0.99 and an initial learning rate of 10-4. Early stopping is enabled after 100 epochs using a patience of 20 epochs and a minimum loss improvement of 10-5. The generalizable INR was trained for up to three days on a single NVIDIA GeForce GTX 1080 Ti.
@@ 3.3 Registration and Atlas Generation
The learnable atlas is a key component to enable resolution independence. It builds upon the idea proposed by Großbröhmer et al. (2024) of representing the atlas using an MLP with adjustable parameters . As such, the atlas is trained using warped input coordinates to predict the most similar subject representation. In this work, we extend this approach by incorporating a combined representation of intensity values and segmentation labels, similar to our previously proposed interpolation method. Thus, we introduce an INR , which takes warped 3D coordinates to reconstruct the intensity value of the OCT image and the class label of the retinal layers. Since we aim to construct a single atlas volume that represents the whole dataset, this INR is (different from our previous approach) non-generalizable. This continuous representation of the atlas enables the intensity and segmentation values to be sampled instantly, without the need for explicit interpolation in the registration process.
The displacement model is a generalizable INR, denoted as with trainable parameters . In contrast to the proposed method for interpolation, the displacement model uses a hypernetwork to modulate its interim outputs, similar to Dannecker et al. (2024) and Dupont et al. (2022). To do so, we use the hypernetwork (depicted in purple in Fig. 3) – consisting of one linear layer for each layer of the INR (depicted in turquoise in Fig. 3) – to predict individual scaling and shifting vectors based on the latent prior input . Let be the hidden size of the INR. The individual scaling vector and shifting vector are the result of the linear mapping with weight matrix , and bias . By modulating with scaling vector and shifting vector , the output of a linear layer with input , weight matrix , bias and activation function becomes .
The resulting latent code-driven architecture is then used to represent a continuous displacement field for each subject as already introduced for classic INRs by Wolterink et al. (2022). Thus, this model gets the original 3D input coordinates and a latent prior . For each coordinate , we thereby obtain a subject-specific displacement
@@
In contrast to the architecture of Dannecker et al. (2024), we separate the parts and of the latent prior that predict and and use two normal distributions and to initialize them. With and the resulting generalizable INR corresponds to a classic INR. This initialization therefore reduces the disturbance from a non-trained latent prior right after initialization.
Training and inference for atlas generation work analogously to those we used for B-scan interpolation. Assembled into one framework, a subset of coordinates of an OCT volume is used to compute the reconstruction and alignment error as
@@ 3.3.1 Implementation Details
The atlas representation is a four-layer MLP with two additional separate layers for reconstruction and segmentation . Each hidden layer has 256 nodes and uses a sinusoidal activation function as proposed by Sitzmann et al. (2020). The prediction heads are based on sigmoid activation for the image intensity values and softmax activation to predict the class values as one-hot encoding. Different from the interpolation task, here, the class labels are redefined to ten retinal layers by merging very thin layers and discarding the choroid label. While Großbröhmer et al. (2024) suggest using a warm-start procedure by increasing the learning rate, we propose to pre-train the atlas INR for 500 epochs to represent the voxel-wise median of the dataset, resulting in a less blurred initialization. The generalizable INR used to model the deformation consists of four layers with 128 nodes and uses sinusoidal activations throughout, including a single sine activation in the final layer to predict the three-dimensional displacement vector. The hypernetwork’s layers are three separate linear layers with 256 nodes each. The model weights are initialized by sampling from a uniform distribution for small and smooth initial deformations. The vector length of 128 for the latent space is chosen empirically. All components – the deformation model, the atlas model, and the latent priors – are trained simultaneously for 2,000 iterations using the Adam optimizer with learning rates of 10-5, 10-6, and 10-5, respectively. The optimizer is combined with an exponential learning rate scheduler with a multiplication factor of 0.999. The batch size is set to one subject, with 916,608 coordinates each. This number corresponds to an equidistant downsampling to a sub-volume containing 2316264 voxels. This volume is randomly selected from all shifted variations of downsampled sub-volumes. Training was conducted on two NVIDIA RTX A5000s and completed in approximately eight hours. During inference, one latent prior is trained for each subject for 100 epochs with a learning rate of 510-4 and the same exponential learning rate scheduler as for training.
@@ 4.1 Anisotropic Interpolation
The interpolation capabilities of the proposed INR are first assessed using the healthy dataset. To do so, the network is trained with 80 OCT volumes and SLO images of 40 volunteers using only 16 B-scans per volume as input to the INR, resulting in even higher inter-slice distances than for typical clinical images. The input B-scans are sampled equidistantly from the entire depth of the image volumes, enabling evaluation on intermediate B-scans. For comparison, we use simple linear interpolation, registration-based interpolation (Ehrhardt et al., 2007) and non-generalizable INRs adapted to single OCT images. To segment intermediate B-scans for the interpolation methods, we proceed as follows: For the linear interpolation, we interpolate the label images linearly and round the resulting continuous values to integers. For the registration-based method, we use the deformation fields resulting from the B-scan interpolation and apply them to the label images using nearest neighbor interpolation. The non-generalizable INRs (SingleINRs) are implemented with a slightly changed architecture, since we observed improved performance for the modified architecture. The SingleINR, thus, consists of three linear layers of size 512 with sinusoidal activation functions (Sitzmann et al., 2020). Furthermore, to get the SingleINR to interpolate between the given B-scans instead of overfitting to the given positions, we intentionally use smaller B-scan distances than the real ones so that the input coordinates lie approximately on an isotropic grid.
@@
Since the proposed generalizable INR (GenINR) is the only one of the considered methods that can be used to interpolate and segment unseen test cases, we additionally evaluate the GenINR for the 20 unseen cases from the healthy dataset. For these cases, new latent priors are learned, while the INR weights remain frozen. We repeat this experiment for the pathological data, using 120 OCT and FAF images from 14 patients for training and 26 images from five patients for testing. Here, we use every second B-scan for training, i.e., 13 B-scans per volume, and evaluate performance for the remaining twelve B-scans. Results for the healthy dataset can be found in Tab. 1 and in Tab. 2 for the CSCR dataset. For both datasets, several image similarity metrics, including mean absolute error (MAE), peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM), and learned perceptual image patch similarity (LPIPS), are reported to assess the image reconstruction and interpolation capability of the considered methods. For a consistent evaluation of both proposed approaches, we used the redefined class labels in which the choroid is discarded and very thin layers are merged. To evaluate the segmentation performance, the Dice score averaged over the nine retinal layers, as well as the average symmetric surface distance (ASSD) and Hausdorff distance (HD) of the layer borders, are reported for the healthy dataset. For the CSCR dataset, we evaluate the segmentation performance with Dice scores of the retina () and of the three disease-related fluid accumulations combined into one binary pathology label ().
@@
The quantitative results for the healthy dataset show that our proposed GenINR achieves the best results for three out of four image similarity metrics. Only for the LPIPS metric, the GenINRs underperform, while the simple interpolation methods (linear and registration-based) achieve the highest values. Simple interpolation hardly changes the contrast and noise level, whereas the GenINR interpolation results look comparably smooth, missing the noisy and high-frequency components of the images (cf. Fig. 4), which might explain the LPIPS results. Regarding image segmentation, the proposed method performs better than the instance-based methods, indicating a better interpolation of the retinal shape achieved by generalization using population-based training. Especially in the fovea region, where large shape differences can be observed for neighboring B-scans, the GenINR is the only method that can meaningfully represent the shape, as shown in the examples on the left side in Fig. 4. The SingleINR, in turn, shows a strong overfitting on the slices used for training, hindering it from interpolating the retinal shape correctly. This overfitting also prevents the SingleINR from benefiting from the additional SLO information; in fact, the additional information, which is different at the interpolated positions than at those used for training, causes SingleINR to fail. The GenINR, in turn, is enabled to reconstruct blood vessels, as shown in the examples on the right in Fig. 4. The localization information provided by the en-face SLO enables the correct positioning of blood vessels and the reconstruction of blood vessel shadows in the outer retina. Reconstruction in the fovea region, where fewer blood vessels are located, appears to be more challenging, as no reconstructions of vessels or vessel shadows are notable. Interpolation artifacts from neighboring B-scans in the form of misplaced shadows, as observed in competing methods, do not occur for the GenINR – all observed vessel shadows are correctly placed. Furthermore, the GenINR enables the interpolation and segmentation of unseen test cases, which is not possible with the instance-specific methods. Although a decline in performance for the test cases compared to the training images can be observed, the GenINR still manages to reconstruct the OCT volumes reliably and to represent the retinal layer boundaries with an average ASSD of .
@@
Similar results can be observed for the CSCR data. Again, the GenINR demonstrates better performance than the baseline methods both in terms of image reconstruction and segmentation, except for the LPIPS metric. As for the healthy data, the simple linear interpolation delivers the best results for this metric. As observed before, the retinal shape is best aligned with the generalizable INR (average Dice score of 98.98 % for the retina), and also the pathologies are better interpolated. For the test data, all metrics considered show similar performance to the training data, except for the Dice score of pathologies. Without FAF integration, an average Dice score of 16.84 % is achieved for the interpolated slices of the test images and improved slightly to 17.59 % with the additional FAF modality. Although there is no known clear relationship between FAF and OCT appearance, the INR seems to benefit from the additional modality, indicating a direct correlation between FAF and OCT. The reduced segmentation performance for the test cases indicates that the highly variable appearance of retinal pathologies in CSCR cannot yet be completely covered by the individualization through the latent codes. Some exemplary interpolation results for the CSCR data are shown in Fig. 5. Here, it can be seen that the retinal layers are reconstructed well for B-scans without severe pathological changes. However, no reconstruction of blood vessels can be observed, which we attribute to the distorting effect of pathologies on training. Also, in pathological areas the retinal layers appear more blurred compared to the regions unaffected by fluids. Still, the shape of both the retina and the pathologies is accurately reproduced by the GenINR.
@@
Across all experiments, the runtime for each subject was below one second for linear interpolation, around 19 seconds for registration-based interpolation, approximately three minutes for the GenINR (inference adaptation to a new subject), and about five minutes for the SingleINR.
@@ 4.2 Atlas Registration
This section assesses the usage of generalizable INRs for image registration tasks – evaluating how effectively they capture structural consistency within a population. Registration performance and atlas generation capabilities are evaluated using the healthy dataset and compared with two different approaches. First, we compared our GenINR framework with the learnable implicit neural atlas approach of Großbröhmer et al. (2024) (SINA), which uses B-spline control points for registration. For the purpose of supervised segmentation, SINA is extended by a second INR that acts as a label atlas, which introduces weak supervision into the atlas generation process. In addition, we modified our proposed method with an non-generalizable INR (SingleINR). Since SingleINRs are trained on individual instances, joint optimization of atlas and deformation presents new challenges. Thus, we used the computed voxel-wise median as a fixed atlas that is registered using the inter-subject approach of Wolterink et al. (2022). The SingleINR consists of a SIREN with four hidden linear layers of size 128 (Sitzmann et al., 2020). As a reference, we include the initial alignment to the atlas (GenINR), and the alignment after affine registration (GenINR).
@@
Intensity similarity is evaluated using the SSIM, while anatomical alignment is evaluated using the Dice and the ASSD. Deformation’s integrity and plausibility are evaluated using the percentage of negative Jacobian determinants () and the size of the displacements, as measured by the L1 norm (). All similarity metrics are evaluated on single B-scans and averaged over each volume. The methods are trained using the healthy dataset with 64 equidistantly sampled B-scans. The experiments are performed using five-fold cross-validation with 80 OCT and segmentation volumes for training and 20 volumes for inference. We conducted four experiments to evaluate the performance of our proposed registration framework. We 1) evaluate the atlas generation capabilities, 2) investigate the registration performance, 3) show that the framework can be used independently of the inference resolution, and 4) perform an ablation study by replacing the implicit atlas representation with an explicit discrete parameter representation.
@@
Our proposed registration framework creates a well-aligned mean representation, i.e. atlas, as indicated by GenINR in Tab. 3. It achieves scores of , , and for ASSD, Dice, and SSIM, respectively. These findings are supported by the qualitative results shown in Fig. 7. We consider an atlas to be good when it is sharp – especially on the layer edges, artifact-free, and anatomically plausible. Our framework shows qualitatively sharper and more precise layer separation as well as hardly any artifacts compared to the median representation or SINA. Moreover, SINA exhibits artifacts at the center of the fovea and, in general, a blurrier representation. While the median atlas already shows minimal artifacts and reasonable layer separation, our method achieves higher sharpness and better separation in the upper retinal layers. Although both generated atlases show a consistent alignment between class labels and intensity representations, our method yields more anatomically plausible results, particularly in the innermost retinal layers, with a continuous representation of the RNFL across the retina.
@@
Focusing on the results of the atlas-to-subject registration experiment in Tab. 3, our approach achieved the best results by about , , and for ASSD, Dice, and SSIM, respectively. Significant improvements are shown in Fig. 6 for GenINR over the initial alignment, affine alignment, and the two approaches that were compared: SINA and SingleINR. Meanwhile, the time taken for inference execution is reduced by around four minutes per subject compared to SINA and around one minute per subject compared to the SingleINR – resulting in faster processing at around per subject. However, SINA presents a significant deterioration in the segmentation metrics compared to the base alignment (GenINR), indicating an insufficient number of control points of the interpolated displacement field. SINA and GenINR produce smooth and fold-free deformations without the need for additional regularization. SingleINR exhibits a slight percentage of folding. Fig. 8 presents the anatomical alignment results using the separated ASSD metric for all layers. Due to pre-processing, the bottom layers (RPE, OPR & SVS, and IS/OS) are already well aligned initially and challenge the methods to generate a proper transition from small to larger displacements. Despite this, our framework achieved a uniform performance distributed across all layers, whereas the SingleINR struggled with the upper middle layers GCL, IPL, and INL.
@@
We conducted an ablation study to illustrate the performance gains of the implicit atlas representation. To do so, we replaced the INR with two learnable parameter images for intensities and class labels, which are then transformed using the deformation field and linear interpolation to be compared with the target volume. This explicit atlas representation scales with the size of the three dimensions , which limited the experiment to a resolution of 16 B-scans. The results are shown in Fig. 4 and support the advantage of implicit atlas representation by even improving alignment performance across all segmentation-based metrics. It should be noted that the training time refers to the latent prior training during inference.